Find the energy (in MeV) released when + decay converts 12X23 (atomic mass = 22.994124 u) into 11X23 (atomic mass = 22.989769 u). Notice that the first atomic mass includes the mass of 12 electrons, while the second atomic mass includes the mass of only 11 electrons.
Answer #1

+ decay converts 12X23 (atomic mass = 22.994124 u) into 11X23 (atomic mass = 22.989769 u).

Subtract both masses to find mass defect.

Mass defect=22.994124 -22.989769 u =0.004355 u

1amu = 931.46 MeV

the energy (in MeV) released =0.004355 *931.46 MeV

=4.056 MeV